Metal x react with a non metal y wnd form a compound xy2.mass percentage of x is 63.22%.determine the ratio of Relative atomic mass of x and y


1
Expert's answer
2022-05-17T15:58:03-0400

"w(x)=M(x)\/(M(x)+M(y))=0.6322"

M(x)=0.6322(M(x)+M(y))

M(x)-0.6322M(x)=0.6322 M(y)

0.3678 M(x)=0.6322M(y)

M(x)/M(y)=0.6322/0.3678=1.72